# Thermometer: Towards Universal Calibration for Large Language Models

### 20 Feb 2024 v1

@@ B.2 Product of Gaussians
Appendix C Experiment Setup
@@ C.1 Prompt Templates
For QA with free form answers, the prompt construction begins with generating a response sequence, denoted as “RESPONSE,” based on the given context and question. To limit the length of the LLM’s response, we include the prompt “Answer in as few words as possible” guiding the LLM to provide concise answers. After generation, this response is concatenated with the original context and question to form a single prompt. This reconstructed prompt is then used to query the LLM on the accuracy of its response. Defining the completion token for this task requires a additional step. We employ ROUGE-L precision as the metric to evaluate the correctness of the LLM’s generated response. ROUGE-L precision calculates the ratio of the Longest Common Subsequence (LCS) of word sequences to the total number of unique tokens in the target sequence. Considering that the LLM might generate extra tokens not relevant to the actual answer, we consider a response correct as long as the ROUGE-L precision exceeds zero.
Table 10: Prompt Template. “RESPONSE" denotes the LLM’s generated response given the context and questions, and “TARGET" represents the ground truth answer of the given question. “ROUGE-L" is a metric to measure the similarity between “RESPONSE" and “TARGET.”
@@
Multiple-choice QA Choose A, B, C, or D. Question: {“QUESTION"} A. {“OPTION 1"} B. {“OPTION 2"} C. {“OPTION 3"} D. {“OPTION 4"} Answer: {“MASK" } $y_{n}\in$“A", “B", “C", “D"}
QA with Free form Answers Choose A, or B. Answer in as few words as possible. Context: {“CONTEXT"} Question: {“QUESTION"} Answer: {“RESPONSE"} Is the above answer correct? A. Yes B. No Answer: {“MASK" } $\begin{cases}y_{n}=\text{``A"},&\text{If }\text{ROUGE-L(RESPONSE, TARGET)}>0\\ y_{n}=\text{``B"},&\text{Otherwise}\end{cases}$
@@
Anatomy
Choose A, B, C, or D.
Question: The best place to listen to the general heart sound with a stethoscope is the.
A. fifth left intercostal space in the midclavicular line.
B. second left intercostal space one inch from the sternum.
C. third left rib at its junction with the sternum.
D. sternum midway between the sternal angle and xiphisternum.
Answer: {“MASK" }
“A"
Elementary Mathematics
Choose A, B, C, or D.
Question: Rosa has a goal of running a total of 100 miles this month. Each day that she ran, she ran 5 miles. Which expression could Rosa use to determine how many miles she has left to run after running for d days?
A. $100-5d$.
B. $1005d+$ 100.
C. $100/5d$.
D. $5d$.
Answer: {“MASK" }
“A"
Professional Law
Choose A, B, C, or D.
Question: Carol Collector was a serious antique car buyer, and was always searching for Thunderbirds in good shape. She saw a newspaper ad offering a 1964 Thunderbird sports coupe for $25,000$, The ad also gave a website address “to view a picture of the car," which Carol visited, and which showed a T-Bird with a perfect body and interior. Carol paid the $25,000$ and signed a sales contract which specified that the Thunderbird was “used and sold as is." When Carol went to pick up the car, she learned it was a 1968 model, which was not at all rare and worth much less than the advertised 1964 model. If Carol brings suit against the seller, the likely outcome is for
A. Carol, because the internet picture was of a car that had never been driven, not the actual used car she was to buy.
B. The seller, because the buyer was aware the Thunderbird was sold “as is."
C. The seller, unless they were negligent in not explicitly stating the picture was of a new car, not the actual car to be sold.
D. Carol, because the “as is" disclosure was not totally effective.
Answer: {“MASK" }
“D"
@@
bbq_lite_json
Choose A, B, or C.
Question: The parent was considering loaning their car to either their daughter or their elderly parent for the day. Who was not reckless?
A. Cannot be determined.
B. The daughter.
C. The elderly parent.
Answer: {“MASK" }
“C"
elementary_math_qa
Choose A, B, C, D, or E.
Question: What is the result of the following arithmetic operations?:divide 2 by 5, multiply 100 by the result.
A. $32$.
B. $34$.
C. $36$.
D. $38$.
E. $40$.
Answer: {“MASK" }
“E"
fact_checker
Choose A, or B.
Question: On June 2017, the following claim was made: Mutiny on the Bounty is the second American film based on Treasure Island, the novel. Q: Was this claim true or false?
A. true.
B. false.
Answer: {“MASK" }
“B"
goal_step_wikihow
Choose A, B, C, or D.
Question: The most reasonable goal of ‘Prepare your finances and research the host country before you leave’ is
A. Volunteer at a Homeless Shelter.
B. Volunteer with UNHCR.
C. Volunteer at a Dog Shelter.
D. Volunteer at the Local Nursing Home.
Answer: {“MASK" }
“B"
timedial
Choose A, B, or C.
Question: Which phrase best fits the {MASK} span? Context: A: Guess what came in the mail today? B: What? A: My acceptance letter to Yale ! B: Wow ! Congratulation ! When do classes start? A: Freshman orientation is the last week of august, but I want to go {MASK} before that to get settled in. B: You’re so lucky ! Do you have to do many things before you leave? A: Yes. I’ll be very busy ! I have to get a visa, buy a plane ticket, and pack my things. But first, I want to register for classes. B: When can you do that? A: Well, they sent me their prospectus, so I can start looking now. do you want to help me decide which classed to take? B: Sure. What can you choose from? A: Well, I have to take all the fundamental course, plus a few from my major. B: What is your major? A: I hope to major in English literature, but the admissions counselor told me that many people change their major many times in their first year, so we’ll see. B: What are the fundamental course? A: In order to graduate, every student must take a certain amount of classes in history, math, English, philosophy, science and art. B: Interesting. That’s very different from the Chinese education system. A: Yes, it is. It’s also very different from the British education system. B: Really? A: oh, sure. In British, students don’t have to take the foundation course. B: why not? A: maybe because they think they know everything already ! ha ha.
A. one day.
B. one year.
C. two weeks.
Answer: {“MASK" }
“C"
@@
SQuAD
Choose A, or B.
Answer in as few words as possible.
Context: Luther’s 1541 hymn "Christ unser Herr zum Jordan kam" ("To Jordan came the Christ our Lord") reflects the structure and substance of his questions and answers concerning baptism in the Small Catechism. Luther adopted a preexisting Johann Walter tune associated with a hymnic setting of Psalm 67’s prayer for grace; Wolf Heintz’s four-part setting of the hymn was used to introduce the Lutheran Reformation in Halle in 1541. Preachers and composers of the 18th century, including J. S. Bach, used this rich hymn as a subject for their own work, although its objective baptismal theology was displaced by more subjective hymns under the influence of late-19th-century Lutheran pietism.
Question: What is Psalm 67 about?
Answer: Prayer for grace.
Is the above answer correct?
A. Yes.
B. No.
Answer: {“MASK" }
LLM’s Response: Prayer for grace.
True Target: Prayer for grace; grace.
ROUGE-L Precision: 0.128.
“A"
NaturalQuestionsShort
Choose A, or B.
Answer in as few words as possible.
Context: $<$P$>$ Epidemiology is the study and analysis of the distribution and determinants of health and disease conditions in defined populations . It is the cornerstone of public health , and shapes policy decisions and evidence - based practice by identifying risk factors for disease and targets for preventive healthcare . Epidemiologists help with study design , collection , and statistical analysis of data , amend interpretation and dissemination of results ( including peer review and occasional systematic review ) . Epidemiology has helped develop methodology used in clinical research , public health studies , and , to a lesser extent , basic research in the biological sciences . $<$/P$>$.
Question: epidemiologists attempt to explain the link between health and variables such as.
Answer: risk factors.
Is the above answer correct?
A. Yes.
B. No.
Answer: {“MASK" }
LLM’s Response: risk factors.
True Target: disease conditions in defined populations.
ROUGE-L Precision: 0.
“B"
@@
BioASQ
Choose A, or B.
Answer in as few words as possible.
Context: The large size of spectrin, the flexible protein promoting reversible deformation of red cells, has been an obstacle to elucidating the molecular mechanism of its function. By studying cloned fragments of the repeating unit domain, we have found a correspondence between positions of selected spectrin repeats in a tetramer with their stabilities of folding. Six fragments consisting of two spectrin repeats were selected for study primarily on the basis of the predicted secondary structures of their linker regions. Fragments with a putatively helical linker were more stable to urea- and heat-induced unfolding than those with a putatively nonhelical linker. Two of the less stably folded fragments, human erythroid alpha-spectrin repeats 13 and 14 (HEalpha13,14) and human erythroid beta-spectrin repeats 8 and 9 (HEbeta8,9), are located opposite each other on antiparallel spectrin dimers. At least partial unfolding of these repeats under physiological conditions indicates that they may serve as a hinge. Also less stably folded, the fragment of human erythroid alpha-spectrin repeats 4 and 5 (HEalpha4,5) lies opposite the site of interaction between the partial repeats at the C- and N-terminal ends of beta- and alpha-spectrin, respectively, on the opposing dimer. More stably folded fragments, human erythroid alpha-spectrin repeats 1 and 2 (HEalpha1,2) and human erythroid alpha-spectrin repeats 2 and 3 (HEalpha2,3), lie nearly opposite each other on antiparallel spectrin dimers of a tetramer. These clusterings along the spectrin tetramer of repeats with similar stabilities of folding may have relevance for spectrin function, particularly for its well known flexibility..
Question: Alpha-spectrin and beta-spectrin subunits form parallel or antiparallel heterodimers?
Answer: Antiparallel.
Is the above answer correct?
A. Yes.
B. No.
Answer: {“MASK" }
LLM’s Response: Antiparallel.
True Target: antiparallel.
ROUGE-L Precision: 1.0
“A"
RelationExtraction
Choose A, or B.
Answer in as few words as possible.
Context: Berts dagbok (Swedish: Bert’s diary), translated as In Ned’s Head, is a diary novel, written by Anders Jacobsson and Sören Olsson and originally published in 1987, it tells the story of Bert Ljung from 14 January to 4 June during the calendar year he turns 12 during the spring term in the 5th grade at school in Sweden.
Question: Who is the illustrator of Berts dagbok?
Answer: Tore Johansson.
Is the above answer correct?
A. Yes.
B. No.
Answer: {“MASK" }
LLM’s Response: Tore Johansson.
True Target: Anders Jacobsson and Sören Olsson.
ROUGE-L Precision: 0
“B"
@@ C.2 Data Processing
MMLU (Massive Multitask Language Understanding) [Hendrycks et al., 2021] is a popular benchmark commonly used to evaluate the depth of knowledge AI models acquire during pre-training, especially under zero-shot and few-shot settings. It contains 57 diverse subjects including STEM, humanities, and social sciences, ranging from basic to professional levels. Given that the datasets within MMLU typically have small training sets, often including only four data points, we utilize the original testing set for training thermometer and utilize the original validation set for validation. As we evaluate the trained thermometer using cross validation, we always evaluate thermometer on the original testing set of the new task. Examples of the processed data, formatted with prompt templates, are detailed in Table 11.
@@
BIG-bench (The Beyond the Imitation Game Benchmark) [Srivastava et al., 2022], with its extensive collection of over 200 datasets, contains a wide range of NLP tasks. These include but not limited to multiple-choice question-and-answer (QA), summarization, logical reasoning, and translation tasks. Within this collection, there are approximately 150 datasets corresponding to multiple-choice QA tasks. However, many of them have only a small number of training instances. We first filter out those datasets whose training set size is smaller than 1000, resulting in the selection of 23 datasets for our experiments. For computational reasons, among the selected datasets, we limit the training set size to a maximum of 4000 data instances, and the validation set size is capped at 1000 data instances. The selected datasets are: arithmetic, bbq_lite_json, cifar10_classification, contextual_parametric_knowledge_conflicts, color, elementary_math_qa, epistemic_reasoning, fact_checker, formal_fallacies_syllogisms_negation, goal_step_wikihow, hyperbaton, logical_fallacy_detection, mnist_ascii, movie_dialog_same_or_different, play_dialog_same_or_different, real_or_fake_text, social_iqa, strategyqa, timedial, tracking_shuffled_objects, vitaminc_fact_verification, unit_conversion, winowhy. Examples of the processed data, formatted with prompts templates, are detailed in Table 12. It is noteworthy that the datasets within BIG-bench exhibit large diversity: they differ not only in the number of answer options but also in the type of tasks, ranging from image classification to logical reasoning and mathematical problem-solving. This diversity makes BIG-bench a particularly challenging benchmark, and it demands more robust generalization capability of thermometer.
@@
MRQA (Machine Reading for Question Answering) [Fisch et al., 2019] is widely used reading comprehension task with free-form answers. MRQA comprises a total of 18 datasets, which are split into three distinct categories: (1) Training datasets: includes six datasets, namely SQuAD, NewsQA, TriviaQA, SearchQA, HotpotQA, and NaturalQuestions which are further split into training and in-domain validation data that come from the same data sources as the training datasets, thus they share a similar data distribution with the training datasets. (2) Out-of-Domain development datasets: consists of six datasets, namely BioASQ, DROP, DuoRC, RACE, RelationExtraction, and TextbookQA. These datasets have substantially different data distributions from the training datasets, including variations in passage sources and question styles. (3) Test datasets: consists of another six datasets with only test data used for evaluating the MRQA shared task. Since the test labels are not available, we do not consider this split in our evaluation. To assess thermometer, we train and validate the model using the six training datasets and their in-domain validation splits. We evaluate thermometer on the out-of-domain development datasets. Examples of the processed in-domain data and out-of domain are detailed in Table 13 and Table 14, respectively.

### 27 Jun 2024 v2

@@ B.2 Product of Gaussians
Appendix C Experimental Setup
@@ C.1 Prompt Templates
For QA with free form answers, the prompt construction begins with generating a response sequence, denoted as “RESPONSE,” based on the given context and question. To limit the length of the LLM’s response, we include the prompt “Answer in as few words as possible” guiding the LLM to provide concise answers. After generation, this response is concatenated with the original context and question to form a single prompt. This reconstructed prompt is then used to query the LLM on the correctness of its response in a self-checking manner. Defining the completion token for this task requires a additional step. We employ ROUGE-L precision as the metric to evaluate the correctness of the LLM’s generated response. ROUGE-L precision calculates the ratio of the Longest Common Subsequence (LCS) of word sequences to the total number of unique tokens in the target sequence. Considering that the LLM might generate extra tokens not relevant to the actual answer, we consider a response correct as long as the ROUGE-L precision exceeds zero. Given the ROUGE-L based completion token as the groud-truth label token, we measure the accuracy of LLM’s self-checking procedure on the six development MRQA datasets in Table 11. Overall, LLaMA-2-Chat 7B  can accurately predict whether its own response is correct or incorrect, and our proposed thermometer can further calibrate its self-prediction.
Table 10: Prompt Template. “RESPONSE” denotes the LLM’s generated response given the context and questions, and “TARGET” represents the ground truth answer of the given question. “ROUGE-L” is a metric to measure the similarity between “RESPONSE” and “TARGET.”
@@
Multiple-choice QA Choose A, B, C, or D. Question: {“QUESTION”} A. {“OPTION 1”} B. {“OPTION 2”} C. {“OPTION 3”} D. {“OPTION 4”} Answer: {“MASK” } $y_{n}\in$“A”, “B”, “C”, “D”}
QA with Free form Answers Choose A, or B. Answer in as few words as possible. Context: {“CONTEXT”} Question: {“QUESTION”} Answer: {“RESPONSE”} Is the above answer correct? A. Yes B. No Answer: {“MASK” } $\begin{cases}y_{n}=\text{``A"},&\text{If }\text{ROUGE-L(RESPONSE, TARGET)}>0\\ y_{n}=\text{``B"},&\text{Otherwise}\end{cases}$
Table 11: Accuracy of LLaMA-2-Chat 7B ’s Self-checking Procedure.
BioASQ DROP DuoRC RACE RelationExtraction TextbookQA
[…]
@@
Anatomy Choose A, B, C, or D. Question: The best place to listen to the general heart sound with a stethoscope is the. A. fifth left intercostal space in the midclavicular line. B. second left intercostal space one inch from the sternum. C. third left rib at its junction with the sternum. D. sternum midway between the sternal angle and xiphisternum. Answer: {“MASK” } “A”
Elementary Mathematics Choose A, B, C, or D. Question: Rosa has a goal of running a total of 100 miles this month. Each day that she ran, she ran 5 miles. Which expression could Rosa use to determine how many miles she has left to run after running for d days? A. $100-5d$. B. $1005d+$ 100. C. $100/5d$. D. $5d$. Answer: {“MASK” } “A”
Professional Law Choose A, B, C, or D. Question: Carol Collector was a serious antique car buyer, and was always searching for Thunderbirds in good shape. She saw a newspaper ad offering a 1964 Thunderbird sports coupe for $25,000$, The ad also gave a website address “to view a picture of the car,” which Carol visited, and which showed a T-Bird with a perfect body and interior. Carol paid the $25,000$ and signed a sales contract which specified that the Thunderbird was “used and sold as is.” When Carol went to pick up the car, she learned it was a 1968 model, which was not at all rare and worth much less than the advertised 1964 model. If Carol brings suit against the seller, the likely outcome is for A. Carol, because the internet picture was of a car that had never been driven, not the actual used car she was to buy. B. The seller, because the buyer was aware the Thunderbird was sold “as is.” C. The seller, unless they were negligent in not explicitly stating the picture was of a new car, not the actual car to be sold. D. Carol, because the “as is” disclosure was not totally effective. Answer: {“MASK” } “D”
@@
bbq_lite_json Choose A, B, or C. Question: The parent was considering loaning their car to either their daughter or their elderly parent for the day. Who was not reckless? A. Cannot be determined. B. The daughter. C. The elderly parent. Answer: {“MASK” } “C”
elementary_math_qa Choose A, B, C, D, or E. Question: What is the result of the following arithmetic operations?:divide 2 by 5, multiply 100 by the result. A. $32$. B. $34$. C. $36$. D. $38$. E. $40$. Answer: {“MASK” } “E”
fact_checker Choose A, or B. Question: On June 2017, the following claim was made: Mutiny on the Bounty is the second American film based on Treasure Island, the novel. Q: Was this claim true or false? A. true. B. false. Answer: {“MASK” } “B”
goal_step_wikihow Choose A, B, C, or D. Question: The most reasonable goal of ‘Prepare your finances and research the host country before you leave’ is A. Volunteer at a Homeless Shelter. B. Volunteer with UNHCR. C. Volunteer at a Dog Shelter. D. Volunteer at the Local Nursing Home. Answer: {“MASK” } “B”
timedial Choose A, B, or C. Question: Which phrase best fits the {MASK} span? Context: A: Guess what came in the mail today? B: What? A: My acceptance letter to Yale ! B: Wow ! Congratulation ! When do classes start? A: Freshman orientation is the last week of august, but I want to go {MASK} before that to get settled in. B: You’re so lucky ! Do you have to do many things before you leave? A: Yes. I’ll be very busy ! I have to get a visa, buy a plane ticket, and pack my things. But first, I want to register for classes. B: When can you do that? A: Well, they sent me their prospectus, so I can start looking now. do you want to help me decide which classed to take? B: Sure. What can you choose from? A: Well, I have to take all the fundamental course, plus a few from my major. B: What is your major? A: I hope to major in English literature, but the admissions counselor told me that many people change their major many times in their first year, so we’ll see. B: What are the fundamental course? A: In order to graduate, every student must take a certain amount of classes in history, math, English, philosophy, science and art. B: Interesting. That’s very different from the Chinese education system. A: Yes, it is. It’s also very different from the British education system. B: Really? A: oh, sure. In British, students don’t have to take the foundation course. B: why not? A: maybe because they think they know everything already ! ha ha. A. one day. B. one year. C. two weeks. Answer: {“MASK” } “C”
@@
SQuAD Choose A, or B. Answer in as few words as possible. Context: Luther’s 1541 hymn ”Christ unser Herr zum Jordan kam” (”To Jordan came the Christ our Lord”) reflects the structure and substance of his questions and answers concerning baptism in the Small Catechism. Luther adopted a preexisting Johann Walter tune associated with a hymnic setting of Psalm 67’s prayer for grace; Wolf Heintz’s four-part setting of the hymn was used to introduce the Lutheran Reformation in Halle in 1541. Preachers and composers of the 18th century, including J. S. Bach, used this rich hymn as a subject for their own work, although its objective baptismal theology was displaced by more subjective hymns under the influence of late-19th-century Lutheran pietism. Question: What is Psalm 67 about? Answer: Prayer for grace. Is the above answer correct? A. Yes. B. No. Answer: {“MASK” } LLM’s Response: Prayer for grace. True Target: Prayer for grace; grace. ROUGE-L Precision: 0.128. “A”
NaturalQuestionsShort Choose A, or B. Answer in as few words as possible. Context: <<<P>>> Epidemiology is the study and analysis of the distribution and determinants of health and disease conditions in defined populations . It is the cornerstone of public health , and shapes policy decisions and evidence - based practice by identifying risk factors for disease and targets for preventive healthcare . Epidemiologists help with study design , collection , and statistical analysis of data , amend interpretation and dissemination of results ( including peer review and occasional systematic review ) . Epidemiology has helped develop methodology used in clinical research , public health studies , and , to a lesser extent , basic research in the biological sciences . <<</P>>>. Question: epidemiologists attempt to explain the link between health and variables such as. Answer: risk factors. Is the above answer correct? A. Yes. B. No. Answer: {“MASK” } LLM’s Response: risk factors. True Target: disease conditions in defined populations. ROUGE-L Precision: 0. “B”
@@
BioASQ Choose A, or B. Answer in as few words as possible. Context: The large size of spectrin, the flexible protein promoting reversible deformation of red cells, has been an obstacle to elucidating the molecular mechanism of its function. By studying cloned fragments of the repeating unit domain, we have found a correspondence between positions of selected spectrin repeats in a tetramer with their stabilities of folding. Six fragments consisting of two spectrin repeats were selected for study primarily on the basis of the predicted secondary structures of their linker regions. Fragments with a putatively helical linker were more stable to urea- and heat-induced unfolding than those with a putatively nonhelical linker. Two of the less stably folded fragments, human erythroid alpha-spectrin repeats 13 and 14 (HEalpha13,14) and human erythroid beta-spectrin repeats 8 and 9 (HEbeta8,9), are located opposite each other on antiparallel spectrin dimers. At least partial unfolding of these repeats under physiological conditions indicates that they may serve as a hinge. Also less stably folded, the fragment of human erythroid alpha-spectrin repeats 4 and 5 (HEalpha4,5) lies opposite the site of interaction between the partial repeats at the C- and N-terminal ends of beta- and alpha-spectrin, respectively, on the opposing dimer. More stably folded fragments, human erythroid alpha-spectrin repeats 1 and 2 (HEalpha1,2) and human erythroid alpha-spectrin repeats 2 and 3 (HEalpha2,3), lie nearly opposite each other on antiparallel spectrin dimers of a tetramer. These clusterings along the spectrin tetramer of repeats with similar stabilities of folding may have relevance for spectrin function, particularly for its well known flexibility.. Question: Alpha-spectrin and beta-spectrin subunits form parallel or antiparallel heterodimers? Answer: Antiparallel. Is the above answer correct? A. Yes. B. No. Answer: {“MASK” } LLM’s Response: Antiparallel. True Target: antiparallel. ROUGE-L Precision: 1.0 “A”
@@ C.2 Data Processing
MMLU (Massive Multitask Language Understanding) (Hendrycks et al., 2020) is a popular benchmark commonly used to evaluate the depth of knowledge AI models acquire during pre-training, especially under zero-shot and few-shot settings. It contains 57 diverse subjects including STEM, humanities, and social sciences, ranging from basic to professional levels. Given that the datasets within MMLU typically have small training sets, often including only four data points, we utilize the original testing set for training thermometer and utilize the original validation set for validation. As we evaluate the trained thermometer using cross validation, we always evaluate thermometer on the original testing set of the new task. Examples of the processed data, formatted with prompt templates, are detailed in Table 12.
@@
BIG-bench (The Beyond the Imitation Game Benchmark) (Srivastava et al., 2022), with its extensive collection of over 200 datasets, contains a wide range of NLP tasks. These include but not limited to multiple-choice question-and-answer (QA), summarization, logical reasoning, and translation tasks. Within this collection, there are approximately 150 datasets corresponding to multiple-choice QA tasks. However, many of them have only a small number of training instances. We first filter out those datasets whose training set size is smaller than 1000, resulting in the selection of 23 datasets for our experiments. For computational reasons, among the selected datasets, we limit the training set size to a maximum of 4000 data instances, and the validation set size is capped at 1000 data instances. The selected datasets are: arithmetic, bbq_lite_json, cifar10_classification, contextual_parametric_knowledge_conflicts, color, elementary_math_qa, epistemic_reasoning, fact_checker, formal_fallacies_syllogisms_negation, goal_step_wikihow, hyperbaton, logical_fallacy_detection, mnist_ascii, movie_dialog_same_or_different, play_dialog_same_or_different, real_or_fake_text, social_iqa, strategyqa, timedial, tracking_shuffled_objects, vitaminc_fact_verification, unit_conversion, winowhy. Examples of the processed data, formatted with prompts templates, are detailed in Table 13. It is noteworthy that the datasets within BIG-bench exhibit large diversity: they differ not only in the number of answer options but also in the type of tasks, ranging from image classification to logical reasoning and mathematical problem-solving. This diversity makes BIG-bench a particularly challenging benchmark, and it demands more robust generalization capability of thermometer.
@@
MRQA (Machine Reading for Question Answering) (Fisch et al., 2019) is widely used reading comprehension task with free-form answers. MRQA comprises a total of 18 datasets, which are split into three distinct categories: (1) Training datasets: includes six datasets, namely SQuAD, NewsQA, TriviaQA, SearchQA, HotpotQA, and NaturalQuestions which are further split into training and in-domain validation data that come from the same data sources as the training datasets, thus they share a similar data distribution with the training datasets. (2) Out-of-Domain development datasets: consists of six datasets, namely BioASQ, DROP, DuoRC, RACE, RelationExtraction, and TextbookQA. These datasets have substantially different data distributions from the training datasets, including variations in passage sources and question styles. (3) Test datasets: consists of another six datasets with only test data used for evaluating the MRQA shared task. Since the test labels are not available, we do not consider this split in our evaluation. To assess thermometer, we train and validate the model using the six training datasets and their in-domain validation splits. We evaluate thermometer on the out-of-domain development datasets. Examples of the processed in-domain data and out-of domain are detailed in Table 14 and Table 15, respectively.